at a basketball game, a team made 54 successful shots. they were a combination of 1- and 2- point shots team scored 90 points in all. write and solve a system of equations to find the number of each shot.
Found 2 solutions by ikleyn, mananth:
Answer by ikleyn(52787)

.
at a basketball game, a team made 54 successful shots. they were a combination of 1- and 2- point shots
team scored 90 points in all. write and solve a system of equations to find the number of each shot.
~~~~~~~~~~~~~~~

Let x be the number of the 1-point shots and y be the number of the 2-point shots.

Then

    x +  y = 54    (1)

    x + 2y = 90    (2)

is your system of two equations to find two unknowns x and y.


Subtracting equation (1) from equation (2), you get

         2y - y = 90 - 54

           y    = 36


Then from equation (2), you get

    x = 54 - 36 = 18.


ANSWER.  18 of the 1-point shots and 36 of the 2-point shots.

Solved.



Answer by mananth(16946)

a team made 54 successful shots. they were a combination of 1- and 2- point shots team scored 90 points in all
Let one point shots be x numbers
and 2 point shots be y
x+y =54------------------------(1)
x+2y =90-----------------------(2)
subtract (1) from (2)
y=36
x=18